A motorcycle travelling 24 m/s comes to rest in 8.7 s. What is the force of friction of the road against the bike tires if the bike weighs 8,200 N?


1
Expert's answer
2021-03-02T18:04:52-0500
"F_{fr}\\Delta t=m(v_f-v_i),\n\n\u200b""F_{fr}=\dfrac{m(v_f-v_i)}{\Delta t}=\dfrac{\dfrac{W}{g}(v_f-v_i)}{\Delta t},""F_{fr}=\dfrac{\dfrac{8200\ N}{9.8\ \dfrac{m}{s^2}}\cdot(0-24\ \dfrac{m}{s})}{8.7\ s}=-2308\ N."